OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3ln3 + 3lnc A. ln(27 + c3) B. ln9c3 C. ln27c D. ln27c3

OpenStudy (abb0t):

\[aln(b) = \ln(b^a)\]

OpenStudy (abb0t):

\[\ln(a) + \ln(b) = \ln(ab)\]

OpenStudy (ghazi):

\[3 \ln 3 +3 \ln c =\ln (3)^3+ \ln (c)^3=\ln 27+\ln c^3=\ln(27c^3)\]

OpenStudy (ghazi):

\[\ln a + \ln b = \ln(ab)\]

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:) i am sorry i wasnt supposed to solve the whole thing but i just did it , my bad :(

OpenStudy (anonymous):

its fine thank you i understand it

OpenStudy (ghazi):

:D